# Workshop on Mechanistic Interpretability

### v1

@@ A.1 Gradients Involving Cross-Entropy and the Log-Sum-Exp Function
$d_{Z}\text{LSE}_{i}(Z\theta^{\intercal})$ $=d_{Z}\log\left(1^{\intercal}\exp(Z\theta^{\intercal})\right)=1^{\intercal}\frac{\exp(Z\theta^{\intercal})\odot dZ\,\theta^{\intercal}}{1^{\intercal}\exp(Z\theta^{\intercal})}=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})\odot dZ\,\theta^{\intercal}\right)$
$=\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right):dZ,$
@@
$\sum_{i,k}\sigma_{i}(Z\theta^{\intercal})_{ij}dZ_{ik}\theta_{jk}=\sum_{i,k}\left[\sigma_{i}(Z\theta^{\intercal})_{ij}\theta_{jk}\right]dZ_{ik}=\sum_{j,k}\left[\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right]_{ijk}dZ_{jk}.$
Since the exterior derivative $df(Z)\in\mathbb{R}^{s}$ of a function $f:\mathbb{R}^{s\times f}\to\mathbb{R}^{s}$ has the expression $\nabla f(Z):dZ$ in terms of the gradient $\nabla f(Z)\in\mathbb{R}^{s\times s\times f}$ at $Z$, the conclusion follows. ∎
@@
$\partial_{Z}\left\langle 1,\text{LSE}_{i}(Z\theta^{\intercal})\right\rangle=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right)=\sigma(Z\theta^{\intercal})\theta=\texttt{CA}(Z,\theta).$
@@
$d\,\text{LSE}_{i}(X)=d\log\left(1^{\intercal}\exp(X)\right)=\frac{1^{\intercal}(\exp(X)\odot dX)}{1^{\intercal}\exp(X)}=1^{\intercal}\left(\sigma_{i}(X)\odot dX\right),$
@@
$df(X)$ $=\left\langle 1,d\,\text{LSE}_{i}(X)\right\rangle=\left\langle 1,1^{\intercal}\left(\sigma_{i}(X)\odot dX\right)\right\rangle=\left\langle 11^{\intercal},\sigma_{i}(X)\odot dX\right\rangle$
$=\left\langle\sigma_{i}(X),dX\right\rangle=\left\langle\nabla f(X),dX\right\rangle,$
@@
$d_{Z}\langle C,N(Z,\theta)\rangle=\langle C,\partial_{Z}N(Z,\theta):dZ\rangle=\langle C:\partial_{Z}N(Z,\theta),dZ\rangle,$
@@
$d_{Z}\left\langle 1,\text{LSE}_{i}\left(N(Z,\theta)\right)\right\rangle$ $=\left\langle\sigma_{i}(N(X,\theta)),dN(Z,\theta)\right\rangle=\left\langle\sigma_{i}(N(X,\theta)),\partial_{Z}N(Z,\theta):dZ\right\rangle$
$=\left\langle\sigma_{i}(N(X,\theta)):\partial_{Z}N(Z,\theta),dZ\right\rangle,$
@@
The $Z$-gradient of the cross entropy $L(Z,\theta)$ for a linear model $N(Z,\theta)=Z\theta^{\intercal}$ satisfies $\partial_{Z}L(Z,\theta)=\mathbb{E}\left[\texttt{CA}\left(Z,\theta\right)-C\theta\right].$
@@
$d_{Z}N(Z,\theta)_{ij}=(dZ\,\theta^{\intercal})_{ij}=\sum_{k}dZ_{ik}\theta_{jk}=\sum_{k,l}\delta_{il}\theta_{jk}dZ_{lk}.$
@@
where $\texttt{CA}(Z,\theta)=\sigma_{i}(Z\theta^{\intercal})\theta$ as claimed. ∎
@@
$d_{Z}N(Z,\theta)_{ij}$ $=(dZ\,\theta Z^{\intercal})_{ij}+(Z\theta\,dZ^{\intercal})_{ij}=\sum_{k}dZ_{ik}(\theta Z^{\intercal})_{kj}+\sum_{k}(Z\theta)_{ik}dZ_{jk}$
$=\sum_{k,l}\left(\delta_{il}(\theta Z^{\intercal})_{kj}+\delta_{jl}(Z\theta)_{ik}\right)dZ_{lk}=\sum_{k,l}\left(\delta_{ik}(Z\theta^{\intercal})_{jl}+\delta_{jk}(Z\theta)_{il}\right)dZ_{kl}.$

### v2

@@ Appendix A Supplemental Derivations for Gradient Calculations
$\mathrm{d}_{Z}\text{LSE}_{i}(Z\theta^{\intercal})$ $=\mathrm{d}_{Z}\log\left(1^{\intercal}\exp(Z\theta^{\intercal})\right)=1^{\intercal}\frac{\exp(Z\theta^{\intercal})\odot\mathrm{d}Z\,\theta^{\intercal}}{1^{\intercal}\exp(Z\theta^{\intercal})}=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})\odot\mathrm{d}Z\,\theta^{\intercal}\right)$
$=\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right):\mathrm{d}Z,$
@@
$\sum_{i,k}\sigma_{i}(Z\theta^{\intercal})_{ij}\mathrm{d}Z_{ik}\theta_{jk}=\sum_{i,k}\left[\sigma_{i}(Z\theta^{\intercal})_{ij}\theta_{jk}\right]\mathrm{d}Z_{ik}=\sum_{j,k}\left[\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right]_{ijk}\mathrm{d}Z_{jk}.$
Since the exterior derivative $\mathrm{d}f(Z)\in\mathbb{R}^{s}$ of a function $f:\mathbb{R}^{s\times f}\to\mathbb{R}^{s}$ has the expression $\nabla f(Z):\mathrm{d}Z$ in terms of the gradient $\nabla f(Z)\in\mathbb{R}^{s\times s\times f}$ at $Z$, the conclusion follows. ∎
@@
$\partial_{Z}\left\langle 1,\text{LSE}_{i}(Z\theta^{\intercal})\right\rangle=1^{\intercal}\left(\sigma_{i}(Z\theta^{\intercal})^{\intercal}\operatorname{\bar{\otimes}}\theta\right)=\sigma(Z\theta^{\intercal})\theta=\mathtt{CA}(Z,\theta).$
@@
$\mathrm{d}\,\text{LSE}_{i}(X)=\mathrm{d}\log\left(1^{\intercal}\exp(X)\right)=\frac{1^{\intercal}(\exp(X)\odot\mathrm{d}X)}{1^{\intercal}\exp(X)}=1^{\intercal}\left(\sigma_{i}(X)\odot\mathrm{d}X\right),$
@@
$\mathrm{d}f(X)$ $=\left\langle 1,\mathrm{d}\,\text{LSE}_{i}(X)\right\rangle=\left\langle 1,1^{\intercal}\left(\sigma_{i}(X)\odot\mathrm{d}X\right)\right\rangle=\left\langle 11^{\intercal},\sigma_{i}(X)\odot\mathrm{d}X\right\rangle$
$=\left\langle\sigma_{i}(X),\mathrm{d}X\right\rangle=\left\langle\nabla f(X),\mathrm{d}X\right\rangle,$
@@
$\mathrm{d}_{Z}\langle C,N(Z,\theta)\rangle=\langle C,\partial_{Z}N(Z,\theta):\mathrm{d}Z\rangle=\langle C:\partial_{Z}N(Z,\theta),\mathrm{d}Z\rangle,$
@@
$\mathrm{d}_{Z}\left\langle 1,\text{LSE}_{i}\left(N(Z,\theta)\right)\right\rangle$ $=\left\langle\sigma_{i}(N(X,\theta)),\mathrm{d}N(Z,\theta)\right\rangle=\left\langle\sigma_{i}(N(X,\theta)),\partial_{Z}N(Z,\theta):\mathrm{d}Z\right\rangle$
$=\left\langle\sigma_{i}(N(X,\theta)):\partial_{Z}N(Z,\theta),\mathrm{d}Z\right\rangle,$
@@
The $Z$-gradient of the cross entropy $L(Z,\theta)$ for a linear model $N(Z,\theta)=Z\theta^{\intercal}$ satisfies $\partial_{Z}L(Z,\theta)=\mathbb{E}\left[\mathtt{CA}\left(Z,\theta\right)-C\theta\right].$
@@
$\mathrm{d}_{Z}N(Z,\theta)_{ij}=(\mathrm{d}Z\,\theta^{\intercal})_{ij}=\sum_{k}\mathrm{d}Z_{ik}\theta_{jk}=\sum_{k,l}\delta_{il}\theta_{jk}\mathrm{d}Z_{lk}.$
@@
where $\mathtt{CA}(Z,\theta)=\sigma_{i}(Z\theta^{\intercal})\theta$ as claimed. ∎
@@
$\mathrm{d}_{Z}N(Z,\theta)_{ij}$ $=(\mathrm{d}Z\,\theta Z^{\intercal})_{ij}+(Z\theta\,\mathrm{d}Z^{\intercal})_{ij}=\sum_{k}\mathrm{d}Z_{ik}(\theta Z^{\intercal})_{kj}+\sum_{k}(Z\theta)_{ik}\mathrm{d}Z_{jk}$
$=\sum_{k,l}\left(\delta_{il}(\theta Z^{\intercal})_{kj}+\delta_{jl}(Z\theta)_{ik}\right)\mathrm{d}Z_{lk}=\sum_{k,l}\left(\delta_{ik}(Z\theta^{\intercal})_{jl}+\delta_{jk}(Z\theta)_{il}\right)\mathrm{d}Z_{kl}.$
@@
Appendix B Supplemental Experiment Details
In this section, we describe additional details for the numerical experiment in the main body.
The corrupted Fashion MNIST dataset is split via a 80%-20% train-validation split. For each image, Gaussian noise is then generated anew for each instance when data is loaded. The added Gaussian noise is entrywise independent and identically distributed with mean 0 and standard deviation of $\frac{1}{3}$.
The linear classifier is trained to minimize the cross-entropy loss via an Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 1024 for 100 epochs. We use an Nvidia a100 GPU with 40GB RAM deploying 32 workers with a prefetch factor of 4 for data loading, to expedite the time required for training, which takes less than three minutes to train. A linear classifier on the clean data using e.g. scikit-learn returns an accuracy of around 85%, so the model performance is suitable compared to other off-the-shelf methods for building a linear classifier.